In a certain​ country, the true probability of a baby being a girl is 0.473. Among the next four randomly selected births in the​ country, what is the probability that at least one of them is a boy​?
The probability is ?. Round to 3 decimal
1
Expert's answer
2020-11-10T19:54:42-0500

Probably that among the next four randomly selected births in the​ country baby being a girl is "0.473^4=0.050054665441". In another way will born at least 1 boy, it means among the next four randomly selected births in the​ country probability that at least one of them is a boy equals"1-0.473^4 = 1-0.050054665441 = 0.949945334559 \approx 0.950"